Circular No. 6245

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995ah IN HS 0016+1449 C. C. Popescu, Max-Planck Institut für Astronomie (MPIA), Heidel-

berg; P. Rafanelli, Dipartimento di Astronomia, Padova; U. Hopp, Universitätssternwarte München and MPIA; and H. Elsässer, MPIA, communicate: “During follow-up spectroscopy at the Calar Alto 2.2-m telescope of emission-line galaxy candidates identified on the digitized plates of the Hamburg Quasar Survey (Popescu et al. 1995, A.Ap. Suppl., in press), we obtained two CCD spectra (resolution 1.2 nm, range 380–860 nm) and a red CCD image of an emission-line galaxy (redshift z = 0.0147) that we designate HS 0016+1449 (α = 0^{h}19^{m}10^{s}.9, δ = +15^{o}06′23′′, equinox 2000.0). A spectrogram obtained on 1995 Feb. 2.792 UT is dominated by the features of a type-II supernova close to maximum brightness; the spectrum of the supernova is characterized by broad emission lines with P-Cyg absorption of the Balmer series and He I (587.6 nm) that correspond to a range in expansion velocity of −20 600 to −9000 km/s. The FWZI of Hα is 20 700 km/s, and its centroid is blueshifted by > 1050 km/s with respect to the systemic velocity. Two broad emission blends centered at 455.0 and 524.0 nm in the rest frame are also present. The Balmer lines are visible up to Hγ. We derive magnitudes for the supernova, estimated from its continuum distribution: B = 17.6, V = 17.2. The red image obtained on 1995 Feb. 3.786 shows the supernova at the northern edge of a low-surface-brightness dwarf galaxy with major axis of length 11′′.7 (axial ratio 0.73); SN 1995ah is located 1′′.6 east and 2′′.1 north of the galaxy’s center. The magnitude of the galaxy, as derived from its spectrum, is B = 18.8, corresponding to absolute mag −15.1 (assuming the Hubble constant is 75 km s^{−1} Mpc^{−1}). An earlier spectrogram, obtained on 1994 Oct. 8.141, shows energy distribution and emission-line ratios typical of an H II galaxy where intensive star formation is occurring [L(Hα) > 7 × 10^{40} erg/s], with no sign of the supernova’s spectral signature. Morphology, absolute magnitude, and spectrum classify HS 0016+1449 as a typical blue compact dwarf galaxy.”

R CORONAE BOREALIS J. A. Mattei, AAVSO, reports that this variable has begun to fade for

the first time since Aug. 1993–Jan. 1994, when the star reached mag 8.5 at minimum. Visual magnitude estimates: Oct. 1.03 UT, 6.1 (F. West, Hanover, PA); 6.80, 6.8 (A. Mizser, Budapest, Hungary); 9.13, 7.2 (C. E. Spratt, Victoria, BC); 12.74, 7.9 (M. V. Zanotta, Passo San Marco, Italy); 15.78, 9.1 (Zanotta); 17.72, 10.1 (L. Kiss, Szeged, Hungary).

1995 October 18 Daniel W. E. Green